OpenStudy (calculusxy):

If a large factory sells its new gadgets for $5 each, it can sell 1050 per month, and if it sells the same gadgets for $9, it will sell 900 per month. Assuming the relation between price and sales is linear, predict the monthly sales of gadgets to the nearest whole number if the price is $12.

OpenStudy (calculusxy):

@Nnesha

OpenStudy (hwyl):

CAN YOU SET UP THEIR LINEAR FUNCTIONS?

OpenStudy (calculusxy):

I think that the slope will be -75/2 because \(\large \frac{900 - 1050}{9 - 5} = \frac{-150}{4} = \frac{-75}{2}\)

OpenStudy (hwyl):

WHAT ARE YOU DOING?

OpenStudy (calculusxy):

Trying to find out the change in y over change in x

OpenStudy (hwyl):

WHY?

OpenStudy (calculusxy):

Well can you tell me your way?

OpenStudy (hwyl):

THAT'S WHY I ASKED IF YOU CAN SET UP THEIR INDIVIDUAL FUNCTIONS

OpenStudy (calculusxy):

I don't think I can.

OpenStudy (hwyl):

YOU CAN FIND THE SLOPE, BUT YOU NEED TO HAVE A PLAN HOW THAT SLOPE WILL HELP YOU

OpenStudy (calculusxy):

ok